the plane average 1000 kilometers per hour on the first half of the round trip,but heavy winds slowed its speed on the return trip took 600 kilometers per hour. If the entire trip to 6 hours find the total distance.
Answer by mananth(16949)

let the distance be x km
time = d/r
Time forward + time return = total yime
x/1000 + x/600 = 6
LCD = 3000
multiply every term by the LCD
3x+5x=18000
8x= 18000
/8
x= 18000/8
x= 2250 km